The cost data for Evencoat Paint for the year 2019 is as follows:

Month Gallons of
Paint
Produced
Equipment
Maintenance
Expenses
January 110,000 $70,700
February 68,000 66,800
March 71,000 67,000
April 77,000 68,100
May 95,000 69,200
June 101,000 70,300
July 125,000 70,400
August 95,000 68,900
September 95,000 69,500
October 89,000 68,600
November 128,000 72,800
December 122,000 71,450

A. Using the high-low method, express the company’s maintenance costs as an equation where x represents the gallons of paint produced. Then estimate the fixed and variable costs.

Fixed cost $

Variable cost $

B. Predict the maintenance costs if 90,000 gallons of paint are produced.

Maintenance cost $

C. Predict the maintenance costs if 81,000 gallons of paint are produced.

Maintenance cost $
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (72,800-66,800)/(128,000-68,000) = 0.10

Fixed cost = 72,800 - (128,000*0.10) = 60,000

A.

Fixed cost 60,000
Variable cost 0.10

B.

Maintenance cost = (90,000*0.1) + 60,000

= 69,000

C.

Maintenance cost = (81,000*0.1) + 60,000

= 68,100